OpenStudy (anonymous):

find 3 additional points through which the line passes. point (6,-4) m=2/3

OpenStudy (dumbcow):

point-slope form \[y-y1 = m(x-x1)\] \[y+4 = (2/3)(x-6)\] \[y = (2/3)x -8\] plug in 3 x-values to find your 3 points

OpenStudy (anonymous):

so it doesnt matter what the x values are?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

you can also think as follows: slope of 2/3 means right 3 up 2 you have (6,-4) so also (9,-2) "right 3 up 2" and (12,0) and (15,2) and (18,4) and (21,6) and ...